# WARDEN: Multi-Directional Backdoor Watermarks for Embedding-as-a-Service Copyright Protection

### v1

@@ Abstract
Embedding as a Service (EaaS) has become a widely adopted solution, which offers feature extraction capabilities for addressing various downstream tasks in Natural Language Processing (NLP). Prior studies have shown that EaaS can be prone to model extraction attacks; nevertheless, this concern could be mitigated by adding backdoor watermarks to the text embeddings and subsequently verifying the attack models post-publication. Through the analysis of the recent watermarking strategy for EaaS, EmbMarker, we design a novel CSE (Clustering, Selection, Elimination) attack that removes the backdoor watermark while maintaining the high utility of embeddings, indicating that the previous watermarking approach can be breached. In response to this new threat, we propose a new protocol to make the removal of watermarks more challenging by incorporating multiple possible watermark directions. Our defense approach, WARDEN, notably increases the stealthiness of watermarks and empirically has been shown effective against CSE attack. [^1] The code is available at https://github.com/anudeex/WARDEN.git.
@@
[…]
@@ 1 Introduction
Given the recent success of EaaS, the associated vulnerabilities have started to attract attention in security and NLP communities (Xu and He, 2023). As a primary example, model extraction attack, a.k.a. imitation attack, has been proven to be effective in stealing the capability of LLMs (Krishna et al., 2020; Tramèr et al., 2016; He et al., 2021a). To conduct such attacks, the attackers query the victim model and then train their own model based on the collected data. Attackers usually invest far less cost and resources than victims to provide competitive services, as shown in Figure 1.a. Therefore, it is imperative to defend against them, and the most popular tactic is to implant statistical signals (or watermarks) via backdoor techniques.
@@
Beyond intellectual property (IP) infringement, further vulnerabilities have been exposed, such as privacy breaches (He et al., 2022a), more performant surrogate models (Xu et al., 2022), and transferable adversarial attacks (He et al., 2021b). As a result, backdoor watermarks are added to EaaS embeddings enabling post-attack lawsuits because the attack models inherit the watermarks, which could be utilized by EaaS providers to identify them. The first work of this kind uses a pre-determined embedding (vector) as the watermark, which is then incorporated into text embeddings in proportion to trigger words Peng et al. (2023), as illustrated in Figure 1.b. The primary requirements for watermarking methods include: (i) they should not lower the quality of the original application, and (ii) it should be difficult for malicious users to identify or deduce the secret watermark vector (Juuti et al., 2019).
Our first work, CSE attack, challenges the above second point. It involves creating a framework CSE (Clustering, Selection, Elimination) that selects the suspected embeddings with watermarks by comparing the distortion between embedding pairs of the victim model and a benchmark model, then neutralizes the impact of the watermark in the embeddings, as shown in Figure 1.c. Empirical evidence demonstrates that CSE successfully compromises the watermark while preserving high embedding utility. To mitigate the effects of CSE, our second work introduces WARDEN, a multi-directional Watermark Augmentation for Robust DEfeNse mechanism, which uses multiple watermark embeddings to reduce the chance of attackers breaching all of them, as depicted in Figure 1.d. We notice that WARDEN, even with a limited number of watermarks, is successful in countering CSE. Moreover, we design a corresponding verification protocol to allow every watermark the authority to verify copyright violations.
@@ 2.1 Imitation Attacks
Imitation attacks Krishna et al. (2020); Orekondy et al. (2019); Yue et al. (2021); Wallace et al. (2020) duplicate cloud models without access to its internal parameters, architecture, or training data. The attack involves sending queries to the victim model and training a functionally similar surrogate model based on API’s responses (Chandrasekaran et al., 2020; Tramèr et al., 2016). Liu et al. 2022, showed that publicly deployed cloud EaaS APIs are also vulnerable to these attacks. It poses a potential threat to EaaS providers, as attackers can easily extract the deployed model in reduced time and with marginal financial investment. More concerningly, such models can outperform the victim models (Xu et al., 2022). Subsequently, they may release a similar API at a lower cost, thereby violating IP rights and causing harm to the market.
@@ 2.2 Backdoor Attacks and Watermarks
Backdoor attacks (Dai et al., 2019), a significant subcategory of adversarial attacks (Alzantot et al., 2018; Ebrahimi et al., 2018), involves inserting textual triggers into a target model such that the victim model behaves normally until the backdoor is activated. Recent works (Zhang et al., 2023; Chen et al., 2022) have shown that pre-trained LLMs are susceptible to backdoor attacks and transferable to downstream tasks.
Recent research (Li et al., 2022; Tang et al., 2023; Peng et al., 2023) has utilized backdoor as the essential technology to integrate verifiable watermark information in deep learning models. The reason is that other techniques, such as altering model parameters (Uchida et al., 2017; Lim et al., 2022), need white-box access and are non-transferable in model extraction attacks. Similarly, lexical watermarks He et al. (2022b, c) do not work on embeddings in the EaaS use case. Drawing inspiration from backdoor attacks, one can make EaaS embeddings correspond to a pre-defined watermark when trigger conditions are satisfied. One such work, EmbMarker (Peng et al., 2023), uses just a single embedding and adds this to original embeddings linearly per the number of moderate-frequency trigger words. However, it was verified against a narrow set of similarity invariant attacks, leaving scope for superior attacks and countermeasures.
@@ 3.1 Preliminary
Malicious attackers target the EaaS victim service $S_{v}$, based on victim model $\Theta_{v}$, by sending texts $t$ to receive corresponding original embeddings ${\bm{e}}_{o}$. Considering the threat of model extraction attacks, the victim backdoors original embedding ${\bm{e}}_{o}$ using a watermarking function $f$ to inject an additional pre-defined embedding $t$ to return provided embedding ${\bm{e}}_{p}=f({\bm{e}}_{o},t)$. Then, the attack model $\Theta_{a}$ is trained on ${\bm{e}}_{p}$ which is received by querying $\Theta_{v}$, and it provides a competitive service $S_{a}$ based on model $\Theta_{a}$. Copyright protection is feasible when $f$ adheres to these criteria: (i) The original EaaS provider should be able to query $S_{a}$ to verify if $\Theta_{a}$ has imitated $\Theta_{v}$; (ii) The utility of provided embeddings ${\bm{e}}_{p}$ is comparable to ${\bm{e}}_{o}$ for downstream tasks.
@@ 3.2 CSE Attack Framework
This section outlines our Clustering, Selection, and Elimination (CSE) attack, as the framework shown in Figure 2. This approach aims at (i) identifying the embedding vectors most likely to contain the watermark and ii) eliminating the influence of the watermark while preserving the essential semantics within the embeddings.
@@
We first employ clustering algorithms to organize the embeddings in the dataset (which attackers have retrieved) into groups. This action enhances the subsequent selection step by: (i) improving the efficiency of calculating pair-wise distance within smaller sets of embeddings, and (ii) providing distinct groups of poisoned data entries, which facilitates the identification of more anomalous pairs. K-Means algorithm (Arthur et al., 2007) is used as the primary clustering approach, while we discuss the effectiveness of other clustering methods in the Appendix C.1. Nevertheless, clustering solely is not sufficient for filtering out the watermarked embeddings. For instance, we can observe from the contour lines in Figure 3 that watermarked samples are spread across clusters and inconspicuous. Furthermore, the centroids of the watermarked samples and overall clusters do not coincide. To counteract this, we thus propose the selection module to identify the most suspicious embeddings with the watermark.
@@
Figure 4: Similarity distribution plot between the target embedding and various embedding types. As we can see, the suspected embeddings returned by the selection module in CSE are distinctly different from unsuspected embeddings and more akin to the target embedding. The results for other datasets can be found in the Appendix B.
@@
We denote the victim model as $\Theta_{v}$ and introduce another hold-out standard model (or benchmark model) as $\Theta_{s}$. Within each cluster $\mathcal{C}_{i}$, we conduct pairwise evaluations on the corresponding embeddings ${\bm{e}}_{p}$ (provided embedding) and ${\bm{e}}_{s}$. Those with distinctive distance changes are considered suspected samples.
EmbMarker incorporates varying proportions of a predetermined target embedding into texts containing trigger words. Since the predefined target embedding lacks shared semantic meaning, we hypothesize that the distance between embedding pairs, which have notable contributions from watermarks, exhibit anomaly behavior (validated in  Figure 4) compared to corresponding distances derived from a standard language model, such as BERT (Devlin et al., 2019). Such difference is used to reflect the significance of the abnormal rank of watermarked pairs, as estimated by Equation 1.
$D_{p}=Rank(D_{v})-Rank(D_{s}),$ (1)
where $D_{v}$ and $D_{s}$ are cosine similarity disparities between victim embeddings (i.e., ${\bm{e}}_{p1}$ and ${\bm{e}}_{p2}$) potentially containing watermark, and the standard embeddings  [^3] We use state-of-the-art SBERT (Reimers and Gurevych, 2019) embeddings offering benefits in distance measurement. (i.e., ${\bm{e}}_{s1}$ and ${\bm{e}}_{s2}$) probably without watermarks. $Rank$ function indicates the rank of the similarity score of the pair in the set of scores. The distinction between the suspected and unsuspected embeddings is illustrated in Figure 4.
@@
Given the suspicious embeddings that are potentially watermarked from the previous step, we hypothesize that the watermark can be identified in suspicious embeddings’ top principal components (validated in Section 4.2) because the target embedding would be common among them. Following this idea, we propose an elimination algorithm, composed of two steps. (i) we apply singular value decomposition (SVD) Golub and Reinsch (1970) to analyze and identify the top $K$ principal components. (ii) Then, the contribution of these top $K$ components are iteratively eliminated using Gram-Schmidt (GS) process  Trefethen and Bau (1997). The elimination step for each principal component vector is demonstrated as follows:
@@
Here, ${\bm{e}}^{\langle 0\rangle}$ is initialized with ${\bm{e}}_{p}$. After each iteration, ${\bm{e}}^{\langle k+1\rangle}$ is acquired by normalizing ${\bm{u}}^{\langle k+1\rangle}$ such that $\text{Norm}\left({\bm{u}}^{\langle k+1\rangle}\right)=\frac{{\bm{u}}^{\langle k+1\rangle}}{||{\bm{u}}^{\langle k+1\rangle}||}$.
@@ 3.3 WARDEN Defense Framework
To diversify the possibility of watermark directions (or target embeddings), we introduce multiple watermarks, noted as ${\bm{W}}=\{{\bm{w}}_{1},{\bm{w}}_{2},...,{\bm{w}}_{R}\}$. This strategy increases the difficulty of inferring all of them via the elimination module in CSE attack. These watermarks remain confidential on servers and can be subject to regular updates. We randomly split the trigger words set, $T$ into $R$ independent subsets $T_{r}$ for $R$ watermarks. Then the trigger counting function, $\lambda_{r}$ is the frequency of trigger words in $T_{r}$ set with a maximum threshold of $m$ (level of watermark). Finally, we add watermarks to the original embedding ${\bm{e}}_{o}$ for text $S$ to generate the provided embedding ${\bm{e}}_{p}$ as follows:
${\bm{e}}_{p}=\text{Norm}\left((1-\sum_{r=1}^{R}\lambda_{r}(S))\cdot{\bm{e}}_{o}+\sum_{r=1}^{R}\lambda_{r}(S)\cdot{\bm{w}}_{r}\right).$ (4)
One thing to note is that because we split $T$ for multiple watermarks, the proportion of watermarked samples is independent of $R$ and is the same as in the single watermark case. It is due to weight values implicitly normalised $\lambda(S)=\lambda_{1}(S)+\lambda_{2}(S)+…+\lambda_{R}(S)$, where $\lambda(S)$ is single watermark weight.
@@
We adopt a conservative approach to copyright verification with multiple watermarks, i.e., if any watermark confidently flags IP infringement, we consider it positive. Hence, we build verification datasets, backdoor texts $D_{b}$ and benign text $D_{n}$ as follows:
$D_{b_{j}}$ =$=\{[t_{1},t_{2},...,t_{m}]|t_{i}\in T_{j}\},\forall j\in[1..R],$ (5)
@@
The premise is that embeddings for these backdoor texts will be closer to their corresponding target embedding in contrast to benign texts in the case of watermarks. We leverage this behavior of embedding backdoors to verify copyright infringement at each watermark level. We quantify the closeness by computing cosine similarity and squared $L_{2}$ distance between target embeddings ${\bm{W}}$ and embeddings of $D_{b}$ and $D_{n}$.
$\cos_{ij}$ $=\frac{{\bm{e}}_{i}\cdot{\bm{w}}_{j}}{||{\bm{e}}_{i}||\cdot||{\bm{w}}_{j}||},\quad l_{2ij}=\biggl{|}\biggl{|}\frac{{\bm{e}}_{i}}{||{\bm{e}}_{i}||}-\frac{{\bm{w}}_{j}}{||{\bm{w}}_{j}||}\biggl{|}\biggl{|}^{2},$ (6)
$C_{b_{j}}$ =$=\{\cos_{ij}|i\in D_{b_{j}}\},C_{n_{j}}=\{\cos_{ij}|i\in D_{n}\},$
$L_{b_{j}}$ =$=\{l_{2ij}|i\in D_{b_{j}}\},L_{n_{j}}=\{l_{2ij}|i\in D_{n}\},$
where $j\in[1..R]$.
@@
$\Delta_{\cos_{k}}$ $=\frac{1}{|C_{b_{k}}|}\sum_{i\in C_{b_{k}}}i-\frac{1}{|C_{n_{k}}|}\sum_{j\in C_{n_{k}}}j,$ (7)
$\Delta_{l2_{k}}$ $=\frac{1}{|L_{b_{k}}|}\sum_{i\in L_{b_{k}}}i-\frac{1}{|L_{n_{k}}|}\sum_{j\in L_{n_{k}}}j.$
@@
$\Delta_{\cos}$ $=\max\limits_{1\leq i\leq R}{\Delta_{\cos_{i}}},$ (8)
$\Delta_{l2}$ $=\min\limits_{1\leq i\leq R}{\Delta_{l2_{i}}},$
p-value $=\min\limits_{1\leq i\leq R}{\text{p-value}_{i}}.$

### v2

@@ Abstract
Embedding as a Service (EaaS) has become a widely adopted solution, which offers feature extraction capabilities for addressing various downstream tasks in Natural Language Processing (NLP). Prior studies have shown that EaaS can be prone to model extraction attacks; nevertheless, this concern could be mitigated by adding backdoor watermarks to the text embeddings and subsequently verifying the attack models post-publication. Through the analysis of the recent watermarking strategy for EaaS, EmbMarker, we design a novel CSE (Clustering, Selection, Elimination) attack that removes the backdoor watermark while maintaining the high utility of embeddings, indicating that the previous watermarking approach can be breached. In response to this new threat, we propose a new protocol to make the removal of watermarks more challenging by incorporating multiple possible watermark directions. Our defense approach, WARDEN, notably increases the stealthiness of watermarks and has been empirically shown to be effective against CSE attack. [^1] The code is available at https://github.com/anudeex/WARDEN.git.
@@
Anudeex Shetty1∗, Yue Teng1∗, Ke He1, Qiongkai Xu1,2† 1School of Computing and Information System, the University of Melbourne, Australia 2School of Computing, FSE, Macquarie University, Australia {anudeexs,ytten,khhe1}@student.unimelb.edu.au qiongkai.xu@mq.edu.au
@@ 1 Introduction
Given the recent success of EaaS, the associated vulnerabilities have started to attract attention in security and NLP communities (Xu and He, 2023). As a primary example, model extraction attack, a.k.a. imitation attack, has been proven to be effective in stealing the capability of LLMs (Krishna et al., 2020; Tramèr et al., 2016; He et al., 2021a). To conduct such attacks, the attackers query the victim model and then train their own model based on the collected data. Attackers usually invest far less cost and resources than victim to provide competitive services, as shown in Figure 1.a. Therefore, it is imperative to defend against them, and the most popular tactic is to implant statistical signals (or watermarks) via backdoor techniques.
@@
Beyond intellectual property (IP) infringement, further vulnerabilities have been exposed, such as privacy breaches (He et al., 2022a), more performant surrogate models (Xu et al., 2022), and transferable adversarial attacks (He et al., 2021b). As a result, backdoor watermarks are added to EaaS embeddings enabling post-attack lawsuits because the attack models inherit the stealthy watermarks, which could be utilized by EaaS providers to identify them. The first work of this kind uses a pre-determined embedding (vector) as the watermark, which is then incorporated into text embeddings in proportion to trigger words Peng et al. (2023), as illustrated in Figure 1.b. The primary requirements for watermarking methods include: (i) they should not lower the quality of the original application, and (ii) it should be difficult for malicious users to identify or deduce the secret watermark vector (Juuti et al., 2019).
Our first work, CSE attack, challenges the aforementioned second point. It involves creating a framework CSE (Clustering, Selection, Elimination) that selects the suspected embeddings with watermarks by comparing the distortion between embedding pairs of the victim model and a benchmark model, then neutralizes the impact of the watermark on the embeddings, as shown in Figure 1.c. Empirical evidence demonstrates that CSE successfully compromises the watermark while preserving high embedding utility. To mitigate the effects of CSE, our second work introduces WARDEN, a multi-directional Watermark Augmentation for Robust DEfeNse mechanism, which uses multiple watermark embeddings to reduce the chance of attackers breaching all of them, as depicted in Figure 1.d. We notice that WARDEN, even with a limited number of watermarks, is successful in countering CSE. Moreover, we design a corresponding verification protocol to allow every watermark the authority to verify copyright violations.
@@ 2.1 Imitation Attacks
Imitation attacks Krishna et al. (2020); Orekondy et al. (2019); Yue et al. (2021); Wallace et al. (2020) duplicate cloud models without access to its internal parameters, architecture, or training data. The attack involves sending queries to the victim model and training a functionally similar surrogate model based on API’s responses (Chandrasekaran et al., 2020; Tramèr et al., 2016). Liu et al. 2022, showed that publicly deployed cloud EaaS APIs are also vulnerable to these attacks. It poses a potential threat to EaaS providers, as attackers can easily extract the deployed model in reduced time and with marginal financial investment. More concerningly, such models can outperform victim models (Xu et al., 2022) when involving victim model ensemble and domain adaptation. Subsequently, they may release a similar API at a lower cost, thereby violating IP rights and causing harm to the market.
@@ 2.2 Backdoor Attacks and Watermarks
Backdoor attacks (Dai et al., 2019), a significant subcategory of adversarial attacks (Alzantot et al., 2018; Ebrahimi et al., 2018), involves inserting textual triggers into a target model such that the victim model behaves normally until the backdoor is activated. Recent works (Zhang et al., 2023; Chen et al., 2022; Huang et al., 2023) have shown that pre-trained LLMs are susceptible to backdoor attacks and transferable to downstream tasks.
Recent research (Li et al., 2022; Tang et al., 2023; Peng et al., 2023) has utilized backdoor as the essential technology to integrate verifiable watermark information in deep learning models, especially LLMs (Kirchenbauer et al., 2023). The reason is that other techniques, such as altering model parameters (Uchida et al., 2017; Lim et al., 2022), need white-box access and are non-transferable in model extraction attacks. Similarly, lexical watermarks He et al. (2022b, c) do not work on embeddings in the EaaS use case. Drawing inspiration from backdoor attacks, one can correspond EaaS embeddings to a pre-defined watermark when trigger conditions are satisfied. One such work, EmbMarker (Peng et al., 2023), uses just a single embedding and adds this to original embeddings linearly as per the number of moderate-frequency trigger words. However, it was verified against a narrow set of similarity invariant attacks, leaving scope for superior attacks and countermeasures.
@@ 3.1 Preliminary
Malicious attackers target the EaaS victim service $S_{v}$, based on victim model $\Theta_{v}$, by sending texts $t$ as queries to receive corresponding original embeddings ${\bm{e}}_{o}$. Considering the threat of model extraction attacks, the victim backdoors original embedding ${\bm{e}}_{o}$ using a watermarking function $f$ to inject an additional pre-defined embedding $t$ to return provided embedding ${\bm{e}}_{p}=f({\bm{e}}_{o},t)$. Then, the attack model $\Theta_{a}$ is trained on ${\bm{e}}_{p}$ which is received by querying $\Theta_{v}$, and the attacker provides a competitive service $S_{a}$ based on model $\Theta_{a}$. Copyright protection is feasible when $f$ adheres to these criteria: (i) the original EaaS provider should be able to query $S_{a}$ to verify if $\Theta_{a}$ has imitated $\Theta_{v}$; (ii) the utility of provided embeddings ${\bm{e}}_{p}$ is comparable to ${\bm{e}}_{o}$ for downstream tasks.
@@ 3.2 CSE Attack Framework
This section outlines our Clustering, Selection, and Elimination (CSE) attack, as the framework shown in Figure 2. This approach aims at (i) identifying the embedding vectors most likely to contain the watermark and (ii) eliminating the influence of the watermark while preserving the essential semantics within the embeddings.
@@
We first employ clustering algorithms to organize the embeddings in the dataset (which attackers have retrieved) into groups. This action enhances the subsequent selection step by: (i) improving the efficiency of calculating pair-wise distance within smaller sets of embeddings, and (ii) providing distinct groups of poisoned data entries, which facilitates the identification of more anomalous pairs. K-Means algorithm (Arthur et al., 2007) is used as the primary clustering approach, while we discuss the effectiveness of other clustering methods in Appendix C.1. Nevertheless, clustering solely is not sufficient for filtering out the watermarked embeddings. For instance, we can observe from the contour lines in Figure 3 that watermarked samples are spread across clusters and inconspicuous. Furthermore, the centroids of the watermarked samples and overall clusters do not coincide. To counteract this, we thus propose the selection module to identify the most suspicious embeddings with the watermark.
@@
Figure 4: Similarity distribution plot between the target embedding and various embedding types. As we can see, the suspected embeddings returned by the selection module in CSE are distinctly different from unsuspected embeddings and more akin to the target embedding. The results for other datasets are reported in Appendix B.
@@
We denote the victim model as $\Theta_{v}$ and introduce another hold-out standard model (or benchmark model) as $\Theta_{s}$. Within each cluster $\mathcal{C}_{i}$, we conduct pairwise evaluations on the corresponding embeddings ${\bm{e}}_{p}$ (provided embedding) and ${\bm{e}}_{s}$ (standard embedding). Those with distinctive distance changes are considered suspected samples.
EmbMarker incorporates varying proportions of a predetermined target embedding into texts containing trigger words. Since the predefined target embedding lacks shared semantic meaning, we hypothesize that the distance between embedding pairs, which have notable contributions from watermarks, exhibit anomaly behavior (validated in  Figure 4) compared to corresponding distances derived from a standard language model, such as BERT (Devlin et al., 2019). Such difference is used to reflect the significance of the abnormal rank of watermarked pairs, as estimated by
$D_{p}=\text{Rank}(D_{v})-\text{Rank}(D_{s}),$ (1)
where $D_{v}$ and $D_{s}$ are cosine similarity disparities between victim embeddings (i.e., ${\bm{e}}_{p1}$ and ${\bm{e}}_{p2}$) potentially containing watermark, and the standard embeddings  [^3] We use state-of-the-art SBERT (Reimers and Gurevych, 2019) embeddings offering benefits in distance measurement. (i.e., ${\bm{e}}_{s1}$ and ${\bm{e}}_{s2}$) probably without watermarks. The $Rank$ function indicates the rank of the similarity scores by the embedding pairs in the set of scores ($D_{v}$ or $D_{s}$). The distinction between the suspected and unsuspected embeddings is illustrated in Figure 4.
@@
Given the suspicious embeddings that are potentially watermarked from the previous step, we hypothesize that the watermark can be identified and recovered in suspicious embeddings’ top principal components (validated in Section 4.2) because the target embedding would be common among them. Following this idea, we propose an elimination algorithm, composed of two steps. First, we apply singular value decomposition (SVD) Golub and Reinsch (1970) to analyze and identify the top $K$ principal components. Then, the contribution of these components are iteratively eliminated using the Gram-Schmidt (GS) process  Trefethen and Bau (1997). The elimination step for each principal component vector is demonstrated as follows:
@@
Here, ${\bm{e}}^{\langle 0\rangle}$ is initialized with ${\bm{e}}_{p}$. After each iteration, ${\bm{e}}^{\langle k+1\rangle}$ is acquired by normalizing ${\bm{u}}^{\langle k+1\rangle}$ such that $\text{Norm}\left({\bm{u}}^{\langle k+1\rangle}\right)={{\bm{u}}^{\langle k+1\rangle}}/{||{\bm{u}}^{\langle k+1\rangle}||}$.
@@ 3.3 WARDEN Defense Framework
To diversify the possibility of watermark directions (or target embeddings), we introduce multiple watermarks, noted as ${\bm{W}}=\{{\bm{w}}_{1},{\bm{w}}_{2},...,{\bm{w}}_{R}\}$. This strategy increases the difficulty of inferring all of them via the elimination module in CSE attack. These watermarks remain confidential on servers and can be subject to regular updates. We randomly split the trigger words set, $T$ into $R$ independent subsets $T_{r}$ for $R$ watermarks. Then the trigger counting function, $\lambda_{r}$ is the frequency of trigger words in $T_{r}$ set with a maximum threshold of $m$ (level of watermark). Finally, we add watermarks to the original embedding ${\bm{e}}_{o}$ for text $S$ to generate the corresponding embedding ${\bm{e}}_{p}$ as follows:
$\text{Norm}\left((1-\sum_{r=1}^{R}\lambda_{r}(S))\cdot{\bm{e}}_{o}+\sum_{r=1}^{R}\lambda_{r}(S)\cdot{\bm{w}}_{r}\right).$ (4)
One thing to note is that because we split $T$ for multiple watermarks, the proportion of watermarked samples is independent of $R$ and is the same as in the single watermark case. Due to weight values being implicitly normalised $\lambda(S)=\lambda_{1}(S)+\lambda_{2}(S)+…+\lambda_{R}(S)$, where $\lambda(S)$ is watermark weight used on a single trigger set.
@@
We adopt a conservative approach to copyright verification with multiple watermarks, i.e., if any watermark confidently flags IP infringement, we consider it positive. Hence, we build verification datasets, backdoor texts $D_{b_{r}}$ and benign text $D_{n}$ as follows:
$D_{b_{r}}$ =$=\{[t_{1},t_{2},...,t_{m}]|t_{i}\in T_{r}\},\forall r\in[1..R],$ (5)
@@
The premise is that embeddings for these backdoor texts will be closer to their corresponding target embedding in contrast to benign texts in the case of watermarks. We leverage this behavior of embedding backdoors to verify copyright infringement at each watermark level. We quantify the closeness by computing cosine similarity and squared $L_{2}$ distance between target embeddings ${\bm{W}}$ and embeddings of $D_{b_{r}}$ and $D_{n}$, i.e.,
$\cos_{ir}$ $=\frac{{\bm{e}}_{i}\cdot{\bm{w}}_{r}}{||{\bm{e}}_{i}||\cdot||{\bm{w}}_{r}||},\quad l_{2ir}=\biggl{|}\biggl{|}\frac{{\bm{e}}_{i}}{||{\bm{e}}_{i}||}-\frac{{\bm{w}}_{r}}{||{\bm{w}}_{r}||}\biggl{|}\biggl{|}^{2},$ (6)
$C_{b_{r}}$ =$=\{\cos_{ir}|i\in D_{b_{r}}\},C_{n_{r}}=\{\cos_{ir}|i\in D_{n}\},$
$L_{b_{r}}$ =$=\{l_{2ir}|i\in D_{b_{r}}\},L_{n_{r}}=\{l_{2ir}|i\in D_{n}\},$
where $r\in[1..R]$.
@@
$\Delta_{\cos_{r}}$ $=\frac{1}{|C_{b_{r}}|}\sum_{i\in C_{b_{r}}}i-\frac{1}{|C_{n_{r}}|}\sum_{j\in C_{n_{r}}}j,$ (7)
$\Delta_{l2_{r}}$ $=\frac{1}{|L_{b_{r}}|}\sum_{i\in L_{b_{r}}}i-\frac{1}{|L_{n_{r}}|}\sum_{j\in L_{n_{r}}}j.$
@@
$\Delta_{\cos}$ $=\max\limits_{1\leq r\leq R}{\Delta_{\cos_{r}}},$ (8)
$\Delta_{l2}$ $=\min\limits_{1\leq r\leq R}{\Delta_{l2_{r}}},$
p-value $=\min\limits_{1\leq r\leq R}{\text{p-value}_{r}}.$